# CodeEvolve: An open source evolutionary coding agent for algorithm discovery and optimization

### v1

@@ 1 Introduction
The intersection of artificial intelligence (AI) and scientific discovery is entering a new era, marked by a paradigm shift from using AI as a tool for data analysis to using it as automated collaborative agents capable of generating, validating, and experimentally refining novel code solutions. At the forefront of this transformation are Large Language Models, which have demonstrated remarkable capabilities in the generation, understanding, and refinement of complex code (Chen et al., 2021). A framework that leverages the use of LLMs as code-optimization agents can tackle challenging algorithmic problems by exploring the solution space in a way that closely resembles human creativity and iteration (Romera-Paredes et al., 2023).
Google DeepMind’s AlphaEvolve (Novikov et al., 2025) has recently proven that combining LLMs with evolutionary algorithms can yield state-of-the-art solutions to many complex problems, from improving foundational matrix multiplication algorithms to optimizing large-scale data center infrastructure. AlphaEvolve, however, is described in a closed-source whitepaper that provides few technical implementation details. This opacity presents a significant barrier to progress and hinders collective innovation.
@@
  • We perform an empirical evaluation of CodeEvolve on a suite of complex mathematical problems used to benchmark Google DeepMind’s AlphaEvolve (Novikov et al., 2025), demonstrating that our method outperforms the latter on four distinct problems, yielding new state-of-the-art results.
@@ 2 Related Work
Automated generation and optimization of computer programs has long been the domain of Genetic Programming (GP) (Koza, 1992, 1994; Langdon and Poli, 2013), where populations of programs are iteratively improved by operators such as crossover—where multiple programs are combined in order to generate a new one—, and mutation—where a program’s content is altered according to some probabilistic process. Although foundational, classical GP methods often struggle with the semantic complexity of modern programming languages. The recent advent of LLMs represents a paradigm shift. With their demonstrated success in generating high-quality solutions for competitive programming tasks (Li et al., 2022), LLMs have emerged as a powerful engine for program synthesis, capable of serving as sophisticated, semantically aware operators for code improvement.
This synergy has given rise to a new class of techniques, formalized as “Evolution through Large Models” (Lehman et al., 2023; Hemberg et al., 2024). The breakthrough application of this concept was FunSearch (Romera-Paredes et al., 2023), which paired an LLM with a programmatic evaluator to discover novel solutions to open problems in mathematics, proving the viability of the approach for genuine scientific discovery. Building on this idea, Google DeepMind recently unveiled AlphaEvolve (Novikov et al., 2025), a closed-source model that generalizes the FunSearch framework from evolving single mathematical functions to entire codebases and a broader range of algorithmic optimization tasks. This system has been successfully in a wide range of different scenarios: from generating more efficient code for GPU kernels and optimizing the operations of warehouse-scale computers to strengthening results in complexity theory (Nagda et al., 2025). This broad applicability establishes LLM-guided evolution not as a niche technique but as a general-purpose engine for automated problem solving.
@@
Following the interest generated by AlphaEvolve, a few open source and independent research projects have focused on developing LLM-driven evolutionary agents. OpenEvolve (Sharma, 2025) provided an accessible implementation of the core features of LLM-driven evolution, confirming its viability for the wider research community. More recently, ShinkaEvolve (Lange et al., 2025) introduced refinements for sample efficiency, employing a bandit-based LLM ensemble and novelty-based rejection filtering to explore the solution space more effectively. CodeEvolve falls in this class of LLM-driven evolutionary models, and is designed to be broadly applicable to any algorithmic problem where one wishes to optimize one or more quantifiable metrics, while also aiming to foster collaborative and transparent research.
@@
LLMs are notoriously sensitive to prompt variations (Anagnostidis and Bulian, 2024), and a considerable research effort has been directed towards creating systems to automatically design and improve prompts, known in the literature as “meta-prompting” (Suzgun and Kalai, 2024; Zhang et al., 2025). CodeEvolve builds upon evolutionary strategies for improving prompts (Fernando et al., 2023; Chen et al., 2023), mirroring the process being executed for the solutions to the underlying problem. By allowing the LLM to reflect on and rewrite its own instructions, it can unlock more diverse and effective evolutionary pathways.
@@
LLM-driven evolution is part of a broader landscape of AI for scientific discovery, and many distinct approaches have seen major success in recent years. Deep Reinforcement Learning (RL), for instance, has achieved landmark results such as discovering faster matrix multiplication algorithms with AlphaTensor (Fawzi et al., 2022). While incredibly powerful, RL typically requires a more structured environment and a well-defined action space. Other approaches, such as agentic systems, leverage LLMs to reason over scientific hypotheses expressed in natural language (Gottweis et al., 2025). CodeEvolve attempts to find a balance between strictness of the former with the more high-level approach of latter, by simultaneously using LLMs to reason about proposed solutions, while also employing a genetic algorithm for a rigorous exploration of the solution space.
@@ 4 Methodology
Inspired by Google DeepMind’s AlphaEvolve Novikov et al. (2025), CodeEvolve integrates an evolutionary framework with Large Language Models (LLMs) to optimize programs. The architecture is based on the island genetic algorithm (Whitley and Starkweather, 1990). This model maintains multiple populations (islands), each evolving independently and periodically exchanging their best-performing individuals (migration). The migration step occurs according to a predefined graph (migration topology) whose nodes are islands and whose edges determine the migration paths. This approach enhances the amount of concurrent solutions being evaluated and the overall population diversity, while also allowing for successful solutions to propagate across the entire search effort, thus being a core component of our system. At each epoch $t$, every island $i$ maintains a population of prompts $\mathcal{P}_{t}^{i}$ and solutions $\mathcal{S}_{t}^{i}$.
@@ 4.1 LLM Ensemble for Solution Generation
We employ an ensemble of different state-of-the-art LLMs to generate and modify solutions, denoted by LLMEnsemble. Each model in the ensemble is assigned a weight, and for each generation task, a model is sampled according to these weights. Our experiments utilize an ensemble of Google’s GEMINI 2.5 models Comanici et al. (2025), with a higher sampling probability for the faster, more cost-effective model (FLASH) and a lower probability for the more powerful, but expensive (PRO). This strategy balances the high-throughput generation of FLASH with the capacity for significant breakthroughs of PRO Novikov et al. (2025), while also being significantly cheaper, as only using the more expensive model is prohibitive in terms of costs.
@@ 4.2 Evolutionary Operators
    $P(S):=\frac{\text{rk}(S)^{-1}}{\sum_{S^{\prime}\in\mathcal{S}_{t}^{i}}\text{rk}(S^{\prime})^{-1}},$ (2)
@@
4: Sample a solution $S\in\mathcal{S}_{t}^{i}$ (Eq. 2)
5: Sample inspirations $I\subseteq\mathcal{S}^{i}_{t}\setminus\{S\}$ (Eq. 2)
6:if $p<1-p_{\text{explr}}$ then
7:  Collect ancestor solutions $A_{k}(S)$ from $S$ to its root in $\mathcal{S}^{i}_{t}$
8:  $S^{\prime}\leftarrow\texttt{LLMEnsemble}(P(S),A_{k}(S),S,I)$
9:  $P^{\prime}\leftarrow\texttt{NULL}$
10:else
@@ 5 Experiments
In this section, we evaluate CodeEvolve on a set of mathematical problems from the benchmark suite used to validate AlphaEvolve (Novikov et al., 2025). Specifically, we are interested in answering the following questions:
@@ 5.1 Benchmark Problems
for all nonnegative functions $f:\mathbb{R}\mapsto\mathbb{R}$, where $\|\cdot\|_{p}$ denotes the $p$-norm of a given function, and fff*f denotes the convolution operation. Hölder’s inequality immediately yields $C\leq 1$, and mathematicians have been trying to bound $C$ from below by explicitly constructing step functions (Matolcsi and Vinuesa, 2009). We are thus interested in obtaining an algorithm to generate nonnegative step functions that maximizes the ratio between $\|f*f\|_{2}^{2}$ and $\|f*f\|_{1}\|f*f\|_{\infty}$ (c.f. Novikov et al. (2025, Sec. B.2., Appendix B)).
@@
In this problem, we wish to find $n$ $d$-dimensional points in order to minimize the ratio between their maximum to minimum distance. We consider two instances of this problem: $n=16,d=2$, and $n=14,d=3$, referred to as P2.A and P2.B, respectively (c.f. Novikov et al. (2025, Sec. B.8., Appendix B)).
@@
In this problem, we wish to place $n$ circles inside a unit square in order to maximize the sum of their radii. We consider two instances: $n=26$ and $n=32$, referred to as P3.A and P3.B, respectively (c.f. Novikov et al. (2025, Sec. B.12., Appendix B)).
@@
This problem is a relaxation of P3, where we still wish to place $n$ circles in order to maximize the sum of their radii, but now we allow them to be placed in a rectangle of perimeter $4$. We consider the case where $n=21$ (c.f. Novikov et al. (2025, Sec. B.13., Appendix B)).
@@ 6 Discussion
While we have demonstrated state-of-the-art performance on a subset of the AlphaEvolve benchmarks, our work is ongoing. We are actively applying CodeEvolve to the remaining mathematical problems outlined in the AlphaEvolve whitepaper (Novikov et al., 2025, Appendix B). For several problems, such as the Heilbronn problem for triangles and packing hexagons, our preliminary results are already approaching the known state-of-the-art bounds. However, other problems, like improving the lower bound for the Kissing Number problem, appear to be significantly harder and may require further methodological innovations. Our goal is to continue pushing the boundaries of what is achievable with a transparent methodology and a limited computational budget. We believe that future algorithmic enhancements to CodeEvolve will lead to to new breakthroughs.
@@
A significant barrier to progress in LLM-driven algorithm discovery is the closed-source nature of systems like AlphaEvolve (Novikov et al., 2025). This opacity hinders replication, comparative analysis, and collective scientific innovation. Our primary contribution is to address this gap by releasing CodeEvolve as a fully open-source project. By providing the complete framework, experimental configurations, and results, we aim to democratize research in this area. We are a small team with limited resources, and we believe that community collaboration is the most effective path toward accelerating progress. We invite researchers and practitioners to build upon our work, introduce new ideas, and collectively tackle challenging algorithmic problems.
@@ 7 Conclusion & Future Work
For future work, we plan to expand upon the current features by incorporating more sophisticated evolutionary mechanisms, such as promoting increased solution diversity via the MAP-Elites algorithm  (Mouret and Clune, 2015), and a more dynamic scheduling for the exploration-exploitation trade-off via RL methods. We also intend to conduct a large-scale analysis of hyperparameter sensitivity to further optimize performance across different problem domains. Finally, we will apply CodeEvolve to a broader range of scientific and engineering challenges to continue pushing the boundaries of automated algorithm discovery.

### v2

@@ 1 Introduction
The intersection of artificial intelligence (AI) and scientific discovery is entering a new era, marked by a paradigm shift from using AI as a tool for data analysis to using it as automated collaborative agents capable of generating, validating, and experimentally refining novel code solutions. At the forefront of this transformation are Large Language Models, which have demonstrated remarkable capabilities in the generation, understanding, and refinement of complex code (chen2021evaluating). A framework that leverages the use of LLMs as code-optimization agents can tackle challenging algorithmic problems by exploring the solution space in a way that closely resembles human creativity and iteration (funsearch2023).
Google DeepMind’s AlphaEvolve (alphaevolve_whitepaper) has recently proven that combining LLMs with evolutionary algorithms can yield state-of-the-art solutions to many complex problems, from improving foundational matrix multiplication algorithms to optimizing large-scale data center infrastructure. AlphaEvolve, however, is described in a closed-source whitepaper that provides few technical implementation details. This opacity presents a significant barrier to progress and hinders collective innovation.
@@
  • We perform an empirical evaluation of CodeEvolve on a suite of complex mathematical problems used to benchmark Google DeepMind’s AlphaEvolve (alphaevolve_whitepaper), demonstrating that our method outperforms the latter on four distinct problems, yielding new state-of-the-art results.
@@ 2 Related Work
Automated generation and optimization of computer programs has long been the domain of Genetic Programming (GP) (koza1992genetic; Koza1994; langdon2013foundations), where populations of programs are iteratively improved by operators such as crossover—where multiple programs are combined in order to generate a new one—, and mutation—where a program’s content is altered according to some probabilistic process. Although foundational, classical GP methods often struggle with the semantic complexity of modern programming languages. The recent advent of LLMs represents a paradigm shift. With their demonstrated success in generating high-quality solutions for competitive programming tasks (li2022alphacode), LLMs have emerged as a powerful engine for program synthesis, capable of serving as sophisticated, semantically aware operators for code improvement.
This synergy has given rise to a new class of techniques, formalized as “Evolution through Large Models” (lehman2023evolution; Hemberg2024). The breakthrough application of this concept was FunSearch (funsearch2023), which paired an LLM with a programmatic evaluator to discover novel solutions to open problems in mathematics, proving the viability of the approach for genuine scientific discovery. Building on this idea, Google DeepMind recently unveiled AlphaEvolve (alphaevolve_whitepaper), a closed-source model that generalizes the FunSearch framework from evolving single mathematical functions to entire codebases and a broader range of algorithmic optimization tasks. This system has been successfully in a wide range of different scenarios: from generating more efficient code for GPU kernels and optimizing the operations of warehouse-scale computers to strengthening results in complexity theory (nagda2025reinforcedgenerationcombinatorialstructures). This broad applicability establishes LLM-guided evolution not as a niche technique but as a general-purpose engine for automated problem solving.
@@
Following the interest generated by AlphaEvolve, a few open source and independent research projects have focused on developing LLM-driven evolutionary agents. OpenEvolve (openevolve) provided an accessible implementation of the core features of LLM-driven evolution, confirming its viability for the wider research community. More recently, ShinkaEvolve (lange2025shinka) introduced refinements for sample efficiency, employing a bandit-based LLM ensemble and novelty-based rejection filtering to explore the solution space more effectively. CodeEvolve falls in this class of LLM-driven evolutionary models, and is designed to be broadly applicable to any algorithmic problem where one wishes to optimize one or more quantifiable metrics, while also aiming to foster collaborative and transparent research.
@@
LLMs are notoriously sensitive to prompt variations (anagnostidis2024susceptiblellmsinfluenceprompts), and a considerable research effort has been directed towards creating systems to automatically design and improve prompts, known in the literature as “meta-prompting” (suzgun2024metapromptingenhancinglanguagemodels; zhang2025metapromptingaisystems). CodeEvolve builds upon evolutionary strategies for improving prompts (fernando2023promptbreeder; chen2023evoprompting), mirroring the process being executed for the solutions to the underlying problem. By allowing the LLM to reflect on and rewrite its own instructions, it can unlock more diverse and effective evolutionary pathways.
@@
LLM-driven evolution is part of a broader landscape of AI for scientific discovery, and many distinct approaches have seen major success in recent years. Deep Reinforcement Learning (RL), for instance, has achieved landmark results such as discovering faster matrix multiplication algorithms with AlphaTensor (fawzi2022alphatensor). While incredibly powerful, RL typically requires a more structured environment and a well-defined action space. Other approaches, such as agentic systems, leverage LLMs to reason over scientific hypotheses expressed in natural language (gottweis2025ai). CodeEvolve attempts to find a balance between strictness of the former with the more high-level approach of latter, by simultaneously using LLMs to reason about proposed solutions, while also employing a genetic algorithm for a rigorous exploration of the solution space.
@@ 4 Methodology
Inspired by Google DeepMind’s AlphaEvolve alphaevolve_whitepaper, CodeEvolve integrates an evolutionary framework with Large Language Models (LLMs) to optimize programs. The architecture is based on the island genetic algorithm (islands_paper). This model maintains multiple populations (islands), each evolving independently and periodically exchanging their best-performing individuals (migration). The migration step occurs according to a predefined graph (migration topology) whose nodes are islands and whose edges determine the migration paths. This approach enhances the amount of concurrent solutions being evaluated and the overall population diversity, while also allowing for successful solutions to propagate across the entire search effort, thus being a core component of our system. At each epoch $t$, every island $i$ maintains a population of prompts $\mathcal{P}_{t}^{i}$ and solutions $\mathcal{S}_{t}^{i}$.
@@ 4.1 LLM Ensemble for Solution Generation
We employ an ensemble of different state-of-the-art LLMs to generate and modify solutions, denoted by LLMEnsemble. Each model in the ensemble is assigned a weight, and for each generation task, a model is sampled according to these weights. Our experiments utilize an ensemble of Google’s GEMINI 2.5 models gemini, with a higher sampling probability for the faster, more cost-effective model (FLASH) and a lower probability for the more powerful, but expensive (PRO). This strategy balances the high-throughput generation of FLASH with the capacity for significant breakthroughs of PRO alphaevolve_whitepaper, while also being significantly cheaper, as only using the more expensive model is prohibitive in terms of costs.
@@ 4.2 Evolutionary Operators
    $\textrm{Pr}(S):=\frac{\text{rk}(S)^{-1}}{\sum_{S^{\prime}\in\mathcal{S}_{t}^{i}}\text{rk}(S^{\prime})^{-1}},$ (2)
@@
4:if $p<1-p_{\text{explr}}$ then
5:  Sample a solution $S\in\mathcal{S}_{t}^{i}$ and inspirations $I\subseteq\mathcal{S}^{i}_{t}\setminus\{S\}$ according to Eq. 2
6:  Collect ancestor solutions $A_{k}(S)$ from $S$ to its root in $\mathcal{S}^{i}_{t}$
7:  $S^{\prime}\leftarrow\texttt{LLMEnsemble}(P(S),A_{k}(S),S,I)$
8:  $P^{\prime}\leftarrow\texttt{NULL}$
9:else
10:  Sample a solution $S\in\mathcal{S}_{t}^{i}$ and inspirations $I\subseteq\mathcal{S}^{i}_{t}\setminus\{S\}$ uniformly at random
@@ 5 Experiments
In this section, we evaluate CodeEvolve on a set of mathematical problems from the benchmark suite used to validate AlphaEvolve (alphaevolve_whitepaper). Specifically, we are interested in answering the following questions:
@@ 5.1 Benchmark Problems
for all nonnegative functions $f:\mathbb{R}\mapsto\mathbb{R}$, where $\|\cdot\|_{p}$ denotes the $p$-norm of a given function, and fff*f denotes the convolution operation. Hölder’s inequality immediately yields $C\leq 1$, and mathematicians have been trying to bound $C$ from below by explicitly constructing step functions (matolcsi2009improvedboundssupremumautoconvolutions). We are thus interested in obtaining an algorithm to generate nonnegative step functions that maximizes the ratio between $\|f*f\|_{2}^{2}$ and $\|f*f\|_{1}\|f*f\|_{\infty}$ (c.f. alphaevolve_whitepaper).
@@
In this problem, we wish to find $n$ $d$-dimensional points in order to minimize the ratio between their maximum to minimum distance. We consider two instances of this problem: $n=16,d=2$, and $n=14,d=3$, referred to as P2.A and P2.B, respectively (c.f. alphaevolve_whitepaper).
@@
In this problem, we wish to place $n$ circles inside a unit square in order to maximize the sum of their radii. We consider two instances: $n=26$ and $n=32$, referred to as P3.A and P3.B, respectively (c.f. alphaevolve_whitepaper).
@@
This problem is a relaxation of P3, where we still wish to place $n$ circles in order to maximize the sum of their radii, but now we allow them to be placed in a rectangle of perimeter $4$. We consider the case where $n=21$ (c.f. alphaevolve_whitepaper).
@@ 6 Discussion
While we have demonstrated state-of-the-art performance on a subset of the AlphaEvolve benchmarks, our work is ongoing. We are actively applying CodeEvolve to the remaining mathematical problems outlined in the AlphaEvolve whitepaper (alphaevolve_whitepaper, Appendix B). For several problems, such as the Heilbronn problem for triangles and packing hexagons, our preliminary results are already approaching the known state-of-the-art bounds. However, other problems, like improving the lower bound for the Kissing Number problem, appear to be significantly harder and may require further methodological innovations. Our goal is to continue pushing the boundaries of what is achievable with a transparent methodology and a limited computational budget. We believe that future algorithmic enhancements to CodeEvolve will lead to to new breakthroughs.
@@
A significant barrier to progress in LLM-driven algorithm discovery is the closed-source nature of systems like AlphaEvolve (alphaevolve_whitepaper). This opacity hinders replication, comparative analysis, and collective scientific innovation. Our primary contribution is to address this gap by releasing CodeEvolve as a fully open-source project. By providing the complete framework, experimental configurations, and results, we aim to democratize research in this area. We are a small team with limited resources, and we believe that community collaboration is the most effective path toward accelerating progress. We invite researchers and practitioners to build upon our work, introduce new ideas, and collectively tackle challenging algorithmic problems.
@@ 7 Conclusion & Future Work
For future work, we plan to expand upon the current features by incorporating more sophisticated evolutionary mechanisms, such as promoting increased solution diversity via the MAP-Elites algorithm  (mouret2015illuminatingsearchspacesmapping), and a more dynamic scheduling for the exploration-exploitation trade-off via RL methods. We also intend to conduct a large-scale analysis of hyperparameter sensitivity to further optimize performance across different problem domains. Finally, we will apply CodeEvolve to a broader range of scientific and engineering challenges to continue pushing the boundaries of automated algorithm discovery.